# A Mathematical Framework for the Problem of Security for Cognition in Neurotechnology

### v1

@@ Abstract
The rapid advancement in neurotechnology in recent years has created an emerging critical intersection between neurotechnology and security. Implantable devices, non-invasive monitoring, and non-invasive therapies all carry with them the prospect of violating the privacy and autonomy of individuals’ cognition. A growing number of scientists and physicians have made calls to address this issue - which we term Cognitive Security - but applied efforts have been limited. A major barrier hampering scientific and engineering efforts to address Cognitive Security is the lack of a clear means of describing and analyzing relevant problems. In this paper we develop Cognitive Security, a mathematical framework which enables such description and analysis by drawing on methods and results from multiple fields. We demonstrate certain statistical properties which have significant implications for Cognitive Security, and then present descriptions of the algorithmic problems faced by attackers attempting to violate privacy and autonomy, and defenders attempting to obstruct such attempts.
@@ 1 Introduction
Recent years have seen rapid advancements in both brain-computer interfaces (BCI) and technologies for increasingly precise yet non-invasive readout and alteration of brain function. This progress has been rightfully hailed as the beginning of a revolution in the treatment of countless neurological and psychiatric disorders, and the potential for beneficial impact on human lives is immense.[1, 2, 3, 4, 5] Even beyond medical applications, there is talk of eventual mass-market use of such technology, and a number of companies are already applying neurotechnology in limited ways for more mundane consumer uses such as entertainment.
@@
However, these emerging technologies open the door to malicious actors in profound ways, once conceived of only in science fiction [^1] As far back as Anthony Gilmore’s 1932 novel The Affair of the Brains,[6], and perhaps most famously in William Gibson’s novel Neuromancer.[7] yet now rapidly becoming reality. As neuromonitoring and neuromodulation technologies progress, we must begin to consider the near-inevitability that an ability to “read” and “write” (in some sense) on the human operating system will emerge.[8, 1, 2, 3, 4, 5] While the level of detail at which this will eventually be feasible remains to be seen, it is already possible to read out thoughts into speech and text with high accuracy at a nearly normal conversational pace,[4, 5] and there is a great deal of active research in increasing the precision of neuromodulation as well. Historical trends paint a picture of significant risk in the likely outcomes in terms of malicious uses of neurotechnology. Much as the emergence of networking and remote access to computers spurred the development of the mathematical frameworks of Cryptography and Differential Privacy in order to protect data therein,[9, 10, 11, 12, 13] we argue that an analogous framework must be developed for those aspects of security and privacy which are unique to neurotechnology. We term this Cognitive Security. Drawing on recent literature in mathematical modeling in cognitive science and theoretical neuroscience – primarily Hyperdimensional Computing and “Quantum” Cognition – and concepts from a number of other fields, in this paper we present an analog of Differential Privacy for biological cognition.
@@
Hyperdimensional Computing (HDC) refers generally to the field which leverages consequences of what has been termed the “blessing of dimensionality” on representation and computation – particularly in noisy systems.[14, 15, 16] That is, as one considers data vectors of increasingly high dimensions, the information content of these vectors becomes increasingly robust to noise if this information is encoded spatially rather than in more standard forms. These encodings are termed Vector-Symbolic Architectures (VSA) or Holographic Representations, and represent concepts and their relatedness based on degrees of proximity on the surface of some high-dimensional object.[17, 14, 15, 16] Traditionally this is a binary hypercube of many thousands of dimensions, but similar properties hold on hyperspheres as well. This will be discussed further in 2. HC has seen success in both theoretical neuroscience and machine learning – in the latter case particularly in the context of embedded computing.[18, 19]
@@
Biological intelligences – brains, immune systems, and more – tend to have extremely noisy dynamics and yet retain very robust representations and computations. Unlike in an artificial computer, in a biological one there will not necessarily be one precise value corresponding to “apple” and another to “orange”, for example. There are known cases of direct, simple mappings from neurons to information, such as grid cells in spatial navigation. However, these cells have also been argued to work in concert to represent more complex information.[20] Other research has also shown that it is possible in some cases to identify specific neurons which will activate in response to images of only one person.[21, 22] However, even in these cases the simple representations are rendered robust by the high dimensionality of their inputs. In this sense, these small ensembles of cells act as a readout of very high-dimensional data. Additionally, there is good evidence to suggest that neural correlates of many cognitive phenomena require distributed representations.[23, 24] In any case, here we are primarily focused on the phenomenology of cognition, and the results in this paper hold regardless of the specific means by which the brain encodes information biologically. The flexibility of the approach to representation taken in this paper will be elaborated upon in Section 2. In a classical computer “apple” and “orange” would each be associated with some bitstrings, and errors in reading, writing, and other operations could well corrupt the computer’s “understanding” of an object. In other words, a significant fraction of the bits are crucial in aggregate – even with error-correcting codes. This is akin to how genomes work, but moving to brains and other complex biological computers we find there are cases where individual “bits” (say, whether or not one neuron fires within a given time window) are in fact not very important, and to a surprising degree.[15, 16, 23, 24] From the simple observables of human cognition we can see that these neural processes are both computationally sophisticated and based upon a representation which is highly robust.
@@
There is general agreement that the brain does not operate in a manner any more than very loosely analogous to a von Neumann architecture, [^2] Overwhelmingly the most common architecture for classical computers. and some researchers in theoretical neuroscience have argued the brain in fact leverages the statistics of high-dimensional spaces to perform robust computations.[22, 25] In this view, brains operate on an aforementioned “Vector-Symbolic Architecture” (VSA), where vectors in some space encode concepts, with varying degrees of similarity between concepts corresponding to varying degrees of proximity within the space. While VSAs would struggle with noise in a low-dimensional space, as the dimension of a data space increases there is a rapid and nonlinear increase in robustness.
@@
These high-dimensional spaces have been termed hyperspaces by a community of researchers spanning theoretical neuroscience, computational linguistics, and low-power hardware, robotics, and machine learning, and the hypervectors (HV) within them are operated on in what is termed Hyperdimensional Computing (HDC). [^3] In HDC, hyperspace and hypervector are different from the stricter mathematical definition. While the strict definition simply corresponds to N-spaces of $N\geq 4$, in HDC something is considered a hyperspace/hypervector once it approaches or exceeds the order of $10^{3}$. HV are generally taken to mean vectors of dimension roughly on the order of $10^{3}$ or greater, [^4] It is interesting to note that this is a scale already reachable by at least one commercial quantum computer[26, 27, 28, 29] and promised by other manufacturers. Though quantum computing is not the focus of this paper, some of its mathematical results can be translated directly to that field. with increasing dimension offering ever-better performance from a statistical perspective. While there is often discussion in various areas of computer science of the “problem of dimensionality”, HDC leverages an interesting inverse property which has been termed the “blessing of dimensionality”.[30, 31] That is, as the dimension of a space increases, the distribution of distances between random vectors tightens when normalized by the dimension of the space. Specifically, these inter-vector distances cluster ever more strongly around $0.5$ after dividing distances by the space’s dimension. From Brown et al’s original use of the term through the two decades since,[30] an increasing range of applications for this property have been identified.[31, 32, 25, 22, 33, 34, 35, 36, 15, 16, 18, 19, 37, 14] For binary vectors this is an unintuitive yet simple result of the Binomial distribution,[15, 16] and in Section 2 we derive similar statistical results for the mathematical structures of Quantum Computing and “Quantum” Cognition.
@@
“Quantum” Cognition is a model which has gained increasing attention in mathematical psychology and cognitive science in recent decades, and regardless of the perhaps unfortunate choice of name it offers significant advantages over other models of cognition.[38, 39, 40, 41, 42] It cannot be overly stressed that this approach to cognitive science is in no way related to hypotheses asserting connections between quantum mechanics and consciousness. [^5] For discussion of such potential or supposed connections, one can see the Penrose-Hameroff hypothesis, which is generally considered to have been disproven by Derakhshani et al,[43] or the work of science fiction authors such as Greg Egan, Robert J Sawyer, and Mike Chen.[44, 45, 46] Instead, quantum cognition simply takes advantage of the fact that the statistical framework based on projective geometry is able to naturally describe the seeming paradoxes that arise from studying cognition. Devised by von Neumann for quantum mechanics,[47] the projective framework for probability is an alternative to the more traditional set-theoretic probability of Kolmogorov.[48] Humans are known to exhibit a wide range of behaviors and thought processes which would be paradoxical when assessed through the lens of set-based probability, but which turn out to be captured in a remarkably natural way by projective probability.[38, 39] To avoid the potential for confusion or misunderstanding in both this paper and future work on Cognitive Security, we will use the term Projective Cognition rather than “Quantum” Cognition, recognizing that the utility of the modeling approach comes in the form of useful features of complex projective geometry.[38, 39, 40] There are a number of these features which have proven an effective and natural fit for cognitive science, including but not limited to:
@@
  • Contextuality – the fact that different phrasings of questions or different variations on a situation can produce different results. This is captured by the centrality of measurement operators in determining outcomes in complex projective probability.
@@
  • Ordering relations – the phenomenon that the order in which two questions are asked or two events are experienced can alter the probabilities of different responses. This is captured by the non-commutativity of linear algebra.
@@
  • Interference effects – cases where the probabilities of cognitive responses to interacting events violate classical probability. These are captured by superposition.
@@
It should be stressed again that this is not in any way some sort of grand theoretical claim. Rather, it is a simple acknowledgement of three neutral facts. First, that the set-theoretic framework of probabilities in terms of events formulated by Kolmogorov et al – by far the most widely used one – is not the only framework.[47] Second, that the set/event-centric framework fails to capture many key aspects of human cognition.[38, 39] Third, that the projective geometry framework developed by von Neumann et al has, by contrast, proven to be a very natural and effective mathematical description of probability in human cognition.[38, 39] In short, we are simply using a version of statistics which is more effective in the context of cognition. Its historical relationship with quantum mechanics is irrelevant to cognition, just as astrophysics’ use of standard probability does not make it relevant to sociology. This paper’s use of the non-standard term Projective Cognition reflects a desire to make this clear.
@@
With the interdisciplinary stage of the paper set, we will begin by demonstrating the robustness of a holographic approach to quantum computing and Projective Cognition, as it will serve as a core aspect of our proposed framework for Cognitive Security. Inspired by Hyperdimensional Computing and Vector-Symbolic Architectures, we present Projective Holographic Computation as an orthogonal approach to improving the problem of error-tolerance in projective systems like quantum computers (or, for example, AI which could be developed based on Projective Cognition) by leveraging statistical properties of high-dimensional vector spaces. As this paper centers on biological cognition, we focus specifically on one form of Projective Holographic Computation: Projective Holographic Cognition. In the coming sections, we will discuss how it enables analysis of statistical properties of mental states across a range of complexity. The approach is deliberately flexible and robust in its ability to describe not only large numbers of discrete beliefs, preferences, attitudes, etc., but also any number of features of mental states which can be more accurately captured by modeling them as comprised of many individual components.
2 Projective Holographic Cognition

### v4

# bibliography.bib

@@ Abstract
The rapid advancement in neurotechnology in recent years has created an emerging critical intersection between neurotechnology and security. Implantable devices, non-invasive monitoring, and non-invasive therapies all carry with them the prospect of violating the privacy and autonomy of individuals’ cognition. A growing number of scientists and physicians have made calls to address this issue, but applied efforts have been relatively limited. A major barrier hampering scientific and engineering efforts to address these security issues is the lack of a clear means of precisely describing and analyzing relevant problems. In this paper we develop Cognitive Neurosecurity, a mathematical framework which enables such description and analysis by drawing on methods and results from multiple fields. We demonstrate certain statistical properties which have significant implications for Cognitive Neurosecurity, and then describe the algorithmic problems faced by attackers attempting to violate privacy and autonomy, and by defenders attempting to obstruct such attempts.
@@ 1 Introduction
Recent years have seen rapid advancements in both brain-computer interfaces (BCI) and technologies for increasingly precise yet non-invasive readout and alteration of brain function. This progress has been rightfully hailed as the beginning of a revolution in the treatment of countless neurological and psychiatric disorders, and the potential for beneficial impact on human lives is immense.[Alagapan_2023_depression_recovery_dbs, Pandarinath_2017_high_performance_bci_speech_paralyzed_patients, Willett_2021_high_performance_BCI_brain_to_text_communication, Willett_2023_high_performance_speech_neuroprosthesis, Metzger_2023_high_performance_neuroprosthesis_speech_decoding_avatar_control] Even beyond medical applications, there is talk of eventual mass-market use of such technology, and a number of companies are already applying more rudimentary neurotechnology in limited ways for more mundane consumer uses such as entertainment.
@@
These emerging technologies also open the door to malicious actors in profound ways, once conceived of only in science fiction [^1] As far back as Anthony Gilmore’s 1932 novel The Affair of the Brains,[gilmore_2018_affair_of_the_brains], and perhaps most famously in William Gibson’s novel Neuromancer.[Gibson_1984_neuromancer] yet now rapidly becoming reality.[bahr_2011_cyber_universal_access_security_risks, Landau_EEG_based_bci_security] As neuromonitoring and neuromodulation technologies progress, we must begin to consider the near-inevitability that further abilities to “read” and “write” (in some sense) on the human operating system will emerge.[Roelfsema_2018_mind_reading_and_writing_future_of_neurotech, Drew_2023_rise_of_brain_reading_technology, Alagapan_2023_depression_recovery_dbs, Pandarinath_2017_high_performance_bci_speech_paralyzed_patients, Willett_2021_high_performance_BCI_brain_to_text_communication, Willett_2023_high_performance_speech_neuroprosthesis, Metzger_2023_high_performance_neuroprosthesis_speech_decoding_avatar_control] While the level of detail at which this will eventually be feasible remains to be seen, it is already possible to read out thoughts into speech and text with high accuracy at a nearly normal conversational pace,[Willett_2023_high_performance_speech_neuroprosthesis, Metzger_2023_high_performance_neuroprosthesis_speech_decoding_avatar_control] and there is a great deal of active research in increasing the precision of neuromodulation as well. There are also a number of papers which describe experimentally verified means of violating privacy and security and discuss highly plausible risks to autonomy.[bellman_2018_facial_recognition_detection_BCI, Lange_2018_BCI_attack_pin_number_recovery, bernal_2020_brain_implant_neuronal_cyberattacks_experiments_1, bernal_2022_bci_security_review] Historical trends paint a picture of significant risk in the likely outcomes in terms of malicious uses of neurotechnology. Much as the emergence of networking and remote access to computers spurred the development of the mathematical frameworks of Cryptography and Differential Privacy in order to protect data therein,[DuPont_2016_history_of_computer_networks_and_cybersecurity, Dwork_2017_survey_of_reconstruction_attacks_differential_privacy, Vadhan_2017_complexity_of_differential_privacy, Dwork_roth_2014_differential_privacy_textbook, wood_2020_differential_privacy_for_non_technical_audience] we argue that an analogous framework must be developed for those aspects of security and privacy which are unique to neurotechnology.
To our knowledge the first term proposed for this category of problem was ”Neurosecurity”, in a 2009 paper by Denning et al.[Denning_2009_neurosecurity_first_paper_neurosurgery_journal] Suggested terms for violations of such security have included ”brainjacking” [^2] We adopt the term ”brainjacking” given its emphasis of the fact that attacks on neuro-cyber-physical systems are most analogous to the takeover of other cyber-physical systems, such as drones, self-driving cars, robotic systems, and so on. Such cyberattacks are frequently termed ”hijacking”, and so our view is that ”brainjacking” emphasizes the bio-cybernetic – in the general sense first expressed by Wiener – nature of the problem.[wiener_1948_cybernetics_original_book, wiener_1950_cybernetics] from Pycroft et al and ”neural cyberattacks” from Bernal et al.[PYCROFT_2016_brainjacking_bci_security, Bernal_2023_reasons_to_prioritize_BCI_security] A few solutions have been proposed, with the most general thus far likely being the BCI Anonymizer of Bonaci et al, which would protect privacy by filtering raw neurological data before a BCI is allowed to perform any computations.[bonaci_2014_brain_security_exocortex] A number of papers have provided technical methods, either for specific types of BCI or broader approaches meant to be incorporated at a hardware level from the beginning of BCI design.[bernal_2022_bci_security_review, Belkacem_cybersecurity_P300_BCI, Ajrawi_2021_BCI_cybersecurity_RFID_based_framework] These are promising directions in the form of hardware, though to our knowledge all of these methods require extreme care on the part of manufacturers, and are vulnerable to potential BCI analogs of air-gap attacks [^3] Air-gap attacks are ones in which even a computer fully isolated from any networks can still be vulnerable to successful cyberattacks. Methods are highly varied and frequently surprisingly creative. Carrara and Adams provide a broad taxonomic survey of such methods.[Carrara_taxonomy_of_air_gap_attacks] Additionally, a wide variety of potential brainjacking attack types have been described.[bernal_2022_bci_security_review]Prior papers have offered mathematical descriptions of security methods, most notably the work of Bonaci et al which proposes a general description from an information-theoretic perspective.[bonaci_2014_brain_security_exocortex] Because of the precision and fundamental nature of their description from a signals perspective, we construct our definitions of security optimization in Section 4 and Section 5 similarly. Indeed, there are natural interfaces between their work and ours, and their BCI Anonymizer is in some ways similar to the noise injection approach we describe in Section 4.
However, ideally one would like to be able to more directly link cognitive science with the problems of BCI-related security, and theoretical foundations for this have thus far not been provided. Reasons for refining such a link include easier interfacing with psychological experiments, more precise descriptions of the cognitive aspects of BCI security, and the ability to analyze more exotic forms of brainjacking. For example, given BCIs can alter the cognition of individuals, an attacker could potentially avoid detection by individual-centric security approaches by inducing smaller changes in many individuals within a group, lowering their information-theoretic profile at the hardware level and leveraging psychology and sociology to accomplish goals. This is only one example, and we suspect that many types of brainjacking methods targeting multiple individuals and their interactions may be developed in the future. In short, direct hardware security of the various types proposed thus far may not be enough to detect and mitigate all attack types, suggesting that a greater degree of cognitive phenomenology should be incorporated in assessment of various types of attacks.
To this end, in this paper we provide such a cognition-centric approach to BCI security, and term it Cognitive Neurosecurity. This name is chosen over Neurosecurity for two reasons. First, because of our focus on cognition rather than other aspects of brain function. Second, because of the framework’s direct applicability to the generally more policy and national security-focused field which shares the name - though BCI security issues are our focus in this paper. [^4] While the latter will not be addressed in this paper, we are already conducting follow-up work on applying our framework to problems within that field and others. Drawing on recent literature in mathematical modeling in cognitive science and theoretical neuroscience – primarily Hyperdimensional Computing and “Quantum” Cognition – and concepts from a number of other fields, in this paper we present an analog of Differential Privacy for biological cognition. [^5] Ienca et al noted this as a possibility in 2018,[Ienca_2018_brain_leaks_cybersecurity_first_mention_of_differential_privacy], and it has also been briefly discussed at a conceptual level by Bernal et al,[bernal_2022_bci_security_review], but there does not appear to have been any mathematical analysis of the approach to date.
@@
Hyperdimensional Computing (HDC) refers generally to the field which leverages consequences of what has been termed the “blessing of dimensionality” on representation and computation – particularly in noisy systems.[Kanerva_2009_introduction_to_distributed_representations_hyperdimensional_computing, klyeko_survey_hyperdimensional_computing_part_I, klyeko_survey_hyperdimensional_computing_part_II] That is, as one considers data vectors of increasingly high dimensions, the information content of these vectors becomes increasingly robust to noise if this information is encoded spatially rather than in more standard forms. These encodings are termed Vector-Symbolic Architectures (VSA) or Holographic Representations, and represent concepts and their relatedness based on degrees of proximity on the surface of some high-dimensional object.[Kanerva_2009_introduction_to_distributed_representations_hyperdimensional_computing, klyeko_survey_hyperdimensional_computing_part_I, klyeko_survey_hyperdimensional_computing_part_II, Jones_2007_holographic_lexicon_encoding] Traditionally this is a binary hypercube of many thousands of dimensions, but similar properties hold on hyperspheres as well. This will be discussed further in 2.1. HC has seen success in both theoretical neuroscience and machine learning – in the latter case particularly in the context of embedded computing.[Neubert_2019_hyperdimensional_computing_for_robotics, Karunaratne_2020_electronics_with_hyperdimensional_computing]
@@
Biological intelligences – brains, immune systems, and more – tend to have extremely noisy dynamics and yet retain very robust representations and computations. Unlike in an artificial computer, in a biological one there will not necessarily be one precise value corresponding to “apple” and another to “orange”, for example. There are known cases of direct, simple mappings from neurons to information, such as grid cells in spatial navigation. Situations of this sort represent what are termed Localist Representations. However, these cells have also been argued to work in concert to represent more complex information.[moser_place_grid_cells_memory_neurobio] Other research has also shown that it is possible in some cases to identify specific neurons which will activate in response to images of only one person.[quiroga_explicit_encoding_single_neurons, gorban_2019_unreasonable_effectiveness_of_small_neural_ensembles] However, even in these cases the simple representations are rendered robust by the high dimensionality of their inputs. In this sense, these small ensembles of cells act as a readout of very high-dimensional data. Additionally, there is good evidence to suggest that neural correlates of many cognitive phenomena require Distributed Representations.[asim_brain_theory_distributed_and_local_both, Kiefer_conceptual_representations_mind_and_brain] In any case, here we are focused on the phenomenology of cognition in our use of HDC, and remain agnostic to the nature of biological representations so as to improve generalizability of our approach. In other words, the results in this paper hold regardless of the specific means by which the brain encodes information biologically. The flexibility of the approach to representation taken in this paper will be elaborated upon in Section 2.1. In a classical computer “apple” and “orange” would each be associated with some bitstrings, and errors in reading, writing, and other operations could well corrupt the computer’s “understanding” of an object. In other words, a significant fraction of the bits are crucial in aggregate – even with error-correcting codes. This is akin to how genomes work, but moving to brains and other complex biological computers we find there are cases where individual “bits” (say, whether or not one neuron fires within a given time window) are in fact not very important, and to a surprising degree.[klyeko_survey_hyperdimensional_computing_part_I, klyeko_survey_hyperdimensional_computing_part_II, asim_brain_theory_distributed_and_local_both, Kiefer_conceptual_representations_mind_and_brain] From the simple observables of human cognition we can see that these neural processes are both computationally sophisticated and based upon a representation which is highly robust.
@@
There is general agreement that the brain does not operate in a manner any more than very loosely analogous to a von Neumann architecture, [^6] Overwhelmingly the most common architecture for classical computers. and some researchers in theoretical neuroscience have argued the brain in fact leverages the statistics of high-dimensional spaces to perform robust computations.[gorban_2019_unreasonable_effectiveness_of_small_neural_ensembles, Gorban_2020_high_dimensional_brain_blessing_of_dimensionality] In this view, brains operate on an aforementioned “Vector-Symbolic Architecture” (VSA), where vectors in some space encode concepts, with varying degrees of similarity between concepts corresponding to varying degrees of proximity within the space. While VSAs would struggle with noise in a low-dimensional space, as the dimension of a data space increases there is a rapid and nonlinear increase in robustness.
@@
These high-dimensional spaces have been termed hyperspaces by a community of researchers spanning theoretical neuroscience, computational linguistics, and low-power hardware, robotics, and machine learning, and the hypervectors within them are operated on in what is termed Hyperdimensional Computing (HDC). [^7] In HDC, hyperspace and hypervector are different from the stricter mathematical definition. While the strict definition simply corresponds to N-spaces of $N\geq 4$, in HDC something is considered a hyperspace/hypervector once it approaches or exceeds the order of $10^{3}$. hypervectors are generally taken to mean vectors of dimension roughly on the order of $10^{3}$ or greater, [^8] It is interesting to note that this is a scale already reachable by at least one commercial quantum computer[Atom_Computing_2023, Ma_2023_atom_computing_1, Evered_2023_atom_computing_2, Scholl_2023_atom_computing_3] and promised by other manufacturers. Though quantum computing is not the focus of this paper, some of its mathematical results can be translated directly to that field. with increasing dimension offering ever-better performance from a statistical perspective. While there is often discussion in various areas of computer science of the “problem of dimensionality”, HDC leverages an interesting inverse property which has been termed the “blessing of dimensionality”.[brown_1997_original_blessing_of_dimensionality_paper, donoho_summary_paper_blessing_of_dimensionality] That is, as the dimension of a space increases, the distribution of distances between random vectors tightens when normalized by the dimension of the space. Specifically, these inter-vector distances cluster ever more strongly around $0.5$ after dividing distances by the space’s dimension. From Brown et al’s original use of the term through the two decades since,[brown_1997_original_blessing_of_dimensionality_paper] an increasing range of applications for this property have been identified.[li_2018_embracing_blessing_of_dimensionality, pereda_2018_blessing_of_dimensionality_ADHD_EEG, donoho_summary_paper_blessing_of_dimensionality, Gorban_2018_blessing_of_dimensionality_mathematical_foundations, Gorban_2020_high_dimensional_brain_blessing_of_dimensionality, gorban_2019_unreasonable_effectiveness_of_small_neural_ensembles, liu_2016_blessing_of_dimensionality_mixture_data, anderson_2014_blessing_of_dimensionality_large_gaussian_mixtures, klyeko_survey_hyperdimensional_computing_part_I, klyeko_survey_hyperdimensional_computing_part_II, Neubert_2019_hyperdimensional_computing_for_robotics, Karunaratne_2020_electronics_with_hyperdimensional_computing, kim_hyper_dimensional_complex_vector_low-power_communication, Kanerva_2009_introduction_to_distributed_representations_hyperdimensional_computing] For binary vectors this is an unintuitive yet simple result of the Binomial distribution,[klyeko_survey_hyperdimensional_computing_part_I, klyeko_survey_hyperdimensional_computing_part_II] and in Section 2.1 we derive similar statistical results for the mathematical structures of Quantum Computing and “Quantum” Cognition.
@@
“Quantum” Cognition is a model which has gained increasing attention in mathematical psychology and cognitive science in recent decades, and regardless of the perhaps unfortunate choice of name it offers significant advantages over other models of cognition.[Bruza_2015_quantum_cognition_review, Pothos_2022_quantum_cognition_review, Busemeyer_2023_quantum_cognition_models, Yearsley_2016_quantum_cognition_decision_theories, Busemeyer_2015_quantum_cognition_applied_to_psychology, Wang_2013_the_potential_of_quantum_cognition, Wang_2013_quantum_question_order_model, Wang_2014_context_effects_quantum_cognition, Pothos_2009_violations_of_quote_rational_unquote_decisions_quantum_cognition, Trueblood_2012_quantum_cognition_causal_reasoning] It cannot be overly stressed that this approach to cognitive science is in no way related to hypotheses asserting connections between quantum mechanics and consciousness. Instead, quantum cognition simply takes advantage of the fact that the statistical framework based on projective geometry is able to naturally describe the seeming paradoxes that arise from studying cognition. Devised by von Neumann for quantum mechanics,[von_neumann_1955_quantum_probability] the projective framework for probability is an alternative to the more traditional set-theoretic probability of Kolmogorov.[Kolmogorov_1933_foundations_of_probability_theory] Humans are known to exhibit a wide range of behaviors and thought processes which would be paradoxical when assessed through the lens of set-based probability, but which turn out to be captured in a remarkably natural way by projective probability.[Bruza_2015_quantum_cognition_review, Pothos_2022_quantum_cognition_review] To avoid the potential for confusion or misunderstanding in both this paper and future work on Cognitive Neurosecurity, we will use the term Projective Cognition rather than “Quantum” Cognition, recognizing that the utility of the modeling approach comes in the form of useful features of complex projective geometry.[Bruza_2015_quantum_cognition_review, Pothos_2022_quantum_cognition_review, Busemeyer_2023_quantum_cognition_models] There are a number of these features which have proven an effective and natural fit for cognitive science, including but not limited to:
@@
  • Contextuality – the fact that different phrasings of questions or different variations on a situation can produce different results. This is captured by the centrality of measurement operators in determining outcomes in complex projective probability.[Wang_2014_context_effects_quantum_cognition, Bruza_2015_quantum_cognition_review, Pothos_2022_quantum_cognition_review]
@@
  • Ordering relations – the phenomenon that the order in which two questions are asked or two events are experienced can alter the probabilities of different responses. This is captured by the non-commutativity of linear algebra.[Wang_2013_quantum_question_order_model, Wang_2013_the_potential_of_quantum_cognition, Bruza_2015_quantum_cognition_review, Pothos_2022_quantum_cognition_review]
@@
  • Interference effects – cases where the probabilities of cognitive responses to interacting events violate classical probability. These are captured by superposition.[Pothos_2009_violations_of_quote_rational_unquote_decisions_quantum_cognition, Trueblood_2012_quantum_cognition_causal_reasoning, Pothos_2022_quantum_cognition_review]
@@
It should be stressed again that this is not in any way some sort of grand theoretical claim. Rather, it is a simple acknowledgement of three neutral facts. First, that the set-theoretic framework of probabilities in terms of events formulated by Kolmogorov et al – by far the most widely used one – is not the only framework.[von_neumann_1955_quantum_probability] Second, that the set/event-centric framework fails to capture many key aspects of human cognition.[Bruza_2015_quantum_cognition_review, Pothos_2022_quantum_cognition_review] Third, that the projective geometry framework developed by von Neumann et al has, by contrast, proven to be a very natural and effective mathematical description of probability in human cognition.[Bruza_2015_quantum_cognition_review, Pothos_2022_quantum_cognition_review] In short, we are simply using a version of statistics which is more effective in the context of cognition. Its historical relationship with quantum mechanics is irrelevant to cognition, just as astrophysics’ use of standard probability does not make it relevant to sociology. This paper’s use of the non-standard term Projective Cognition reflects a desire to make this clear.
@@
With the interdisciplinary stage of the paper set, we will begin by demonstrating the robustness of a holographic approach to quantum computing and Projective Cognition, as it will serve as a core aspect of our proposed framework for Cognitive Neurosecurity. Inspired by Hyperdimensional Computing and Vector-Symbolic Architectures, we present Projective Holographic Computation as an orthogonal approach to improving the problem of error-tolerance in projective systems like quantum computers (or, for example, AI which could be developed based on Projective Cognition) by leveraging statistical properties of high-dimensional vector spaces. As this paper centers on biological cognition, we focus specifically on one form of Projective Holographic Computation: Projective Holographic Cognition. In the coming sections, we will discuss how it enables analysis of statistical properties of mental states across a range of complexity. The approach is deliberately flexible and robust in its ability to describe not only large numbers of discrete beliefs, preferences, attitudes, etc., but also any number of features of mental states which can be more accurately captured by modeling them as comprised of many individual components.
2 Methods
2.1 Projective Holographic Cognition